# The cohomological Kudla conjecture for unitary Shimura varieties

### v1

@@ 4.1. Poincaré dual forms
For each $\lambda\in M^{\vee}$, let $u^{*}_{\lambda}(\omega_{E})$ denote the pullback of $\omega_{E}$ under the morphism in Equation 3.1. Then $u^{*}_{\lambda}(\omega_{E})\in\mathcal{H}^{1,1}(E_{M},\mathbb{R})$, where $\mathcal{H}^{1,1}(E_{M},\mathbb{R})$ is the real vector space of harmonic $(1,1)$ forms on $E_{M}$ with real coefficients.
@@
$\lambda$ $\mapsto f(\lambda):=u^{*}_{\lambda}(\omega_{E})\leavevmode\nobreak\ .$
@@
$Z(\lambda_{1},\ldots,\lambda_{g})=Z(\lambda_{1})\cap\ldots\cap Z(\lambda_{g})\leavevmode\nobreak\ .$
As for $(3)$, it results from the equality $[a]^{*}\omega_{E}=|a|^{2}\omega_{E}$ valid for $a\in\mathcal{O}_{k}$ and the composition:
$u_{a\lambda}:E_{M}\xrightarrow{h(\lambda,\,\cdot)}E\xrightarrow{[a]}E\leavevmode\nobreak\ .$
@@
$(x_{1},\ldots,x_{n})$ $\mapsto\sum_{\ell=1}^{n}h(\lambda,e_{\ell})\cdot x_{\ell}=\sum_{\ell=1}^{n}a_{\ell}\lambda_{\ell}\cdot x_{\ell}$
@@
$[Y_{\ell}]=a_{\ell}^{2}[E^{\ell-1}\times\{0\}\times E^{n-\ell+1}]\in\mathrm{H}^{2}(E_{M},\mathbb{Z})\leavevmode\nobreak\ .$
@@
$M^{\vee}=\oplus_{\ell=1}^{n}{\mathfrak{D}}^{-1}_{k}\frac{1}{a_{\ell}}e_{\ell}\subset M\otimes_{{\mathcal{O}}_{k}}k\leavevmode\nobreak\ .$
@@
$[Z(\lambda)]=\sum_{\ell=1}^{n}|\lambda_{\ell}|^{2}Y_{\ell}-\sum_{1\leq\ell<j\leq n}\left[\Re(\lambda_{\ell}\overline{\lambda}_{j})Y_{\ell,j}^{+}+\frac{1}{\delta_{k}}\Im(\lambda_{\ell}\overline{\lambda}_{j})Y_{\ell j}^{-}\right]\in\mathrm{H}^{2}(E_{M},\mathbb{Q})\leavevmode\nobreak\ .$
@@
$u_{\lambda}^{*}(dz)=\sum_{\ell=1}^{n}a_{\ell}{\lambda}_{\ell}dz_{\ell}\quad\textrm{and}\quad u_{\lambda}^{*}(d\overline{z})=\sum_{\ell=1}^{n}a_{\ell}\overline{\lambda}_{\ell}d\overline{z}_{\ell}\leavevmode\nobreak\ .$
@@
$Y_{\ell,j}^{+}=\frac{ia_{\ell}a_{j}}{\sqrt{d_{k}}}(dz_{\ell}\wedge d\overline{z}_{j}+dz_{j}\wedge d\overline{z}_{\ell}),\quad\textrm{and}\quad Y_{\ell,j}^{-}=\delta_{k}\frac{ia_{\ell}a_{j}}{\sqrt{d_{k}}}(dz_{\ell}\wedge d\overline{z}_{j}-dz_{j}\wedge d\overline{z}_{\ell})\leavevmode\nobreak\ .$
@@
$[Z(\lambda)]=\sum_{\ell=1}^{n}|\lambda_{\ell}|^{2}Y_{\ell}-\sum_{\ell<j}a_{\ell}a_{j}\Big{(}\Re(\lambda_{\ell}\overline{\lambda}_{j})Y_{\ell,j}^{+}+\Im(\lambda_{\ell}\overline{\lambda}_{j})Y^{-}_{\ell,j}\Big{)}\leavevmode\nobreak\ .$
@@
The class of the divisor $D$ satisfies:
$D=\sum_{\ell=1}^{n}\frac{1}{a_{\ell}}\cdot Y_{\ell}\in{H}^{2}(E_{M},\mathbb{Z})$
@@
The class of $D$ is determined by the Hermitian pairing $h$ on $\mathrm{H}_{1}(E_{M},\mathbb{C})=\mathrm{H}_{1}(E,\mathbb{C})\otimes M=T_{0}E\otimes M\oplus\overline{T_{0}E}\otimes M$. Notice that $D(\partial_{z}\otimes e_{\ell},\partial_{z}\otimes e_{j})=0$ and $D(\partial_{z}\otimes e_{\ell},\partial_{\overline{z}}\otimes e_{j})=\delta_{\ell j}\ \frac{i}{\sqrt{d_{k}}}\cdot a_{\ell}$. Therefore,
$D=\frac{i}{\sqrt{d_{k}}}\sum_{\ell=1}^{n}a_{\ell}\cdot dz_{\ell}\wedge d\overline{z}_{\ell}=\sum_{\ell=1}^{n}\frac{1}{a_{\ell}}\cdot Y_{\ell}\in{H}^{2}(E_{M},\mathbb{Z})\leavevmode\nobreak\ ,$
@@
We have the following relation between $D$ and the Chern class of the conormal bundle of ${\mathcal{B}}_{\mathfrak{J}}$:
$c_{1}({\mathcal{N}}_{{\mathcal{B}}_{{\mathfrak{J}}}}^{\vee})=\frac{d_{k}}{r_{\mathfrak{J}}}\cdot D\leavevmode\nobreak\ .$
@@
This follows from Theorem 4.4 and Equation 3.3. ∎
The following is a consequence of Proposition 4.3, by unicity of harmonic lifts.
@@
$f(\lambda)=\sum_{\ell=1}^{n}|\lambda_{\ell}|^{2}f(e_{\ell})-\sum_{1\leq\ell<j\leq n}[\Re(\lambda_{\ell}\overline{\lambda}_{j})f_{\ell,j}^{+}+\frac{1}{\delta_{k}}\Im(\lambda_{\ell}\overline{\lambda}_{j})f_{\ell,j}^{-}]\in{\mathcal{H}}^{1,1}(E_{M},\mathbb{R})\leavevmode\nobreak\ .$
@@
$f(\lambda_{1},\lambda_{2})=\frac{i}{\sqrt{d_{k}}}u_{\lambda_{1}}^{*}(dz)\wedge u_{\lambda_{2}}^{*}(d\overline{z})\leavevmode\nobreak\ ,$
where $dz$ is the canonical harmonic $1$-from on $E=\mathbb{C}/{\mathcal{O}}_{k}$.
The next proposition shows that $f$ admits an extension to $M_{\mathbb{R}}$. We denote by $\Delta$ be the $\mathrm{SU}(h)$-invariant Laplacian on $M_{\mathbb{R}}$.
@@
$f:M_{\mathbb{R}}\rightarrow\mathcal{H}^{1,1}(E_{M},\mathbb{C})\leavevmode\nobreak\ ,$
which is a Hermitian form and which satisfies $\Delta(f)=D$.
@@
It is clear from Theorem 4.6 that $f$ admits an extension to $M_{\mathbb{R}}$. Moreover, the invariant Laplacian is given by:
$\Delta=\prescript{t}{}{\left(\frac{\partial}{\partial\lambda}\right)}\cdot H^{-1}\cdot\frac{\partial}{\partial\overline{\lambda}}=\sum_{\ell=1}^{n}\frac{1}{a_{\ell}}\frac{\partial^{2}}{\partial\lambda_{\ell}\partial\overline{\lambda}_{\ell}}\leavevmode\nobreak\ ,$
@@
$\Delta(f)(\lambda)=\sum_{\ell=1}^{n}\frac{1}{d_{\ell}}Y_{\ell},$
which proves the desired result by Theorem 4.4. ∎
@@
$\underline{\lambda}$ $\mapsto f(\underline{\lambda})=\det[\left(f(\lambda_{\ell},\lambda_{j})\right)_{1\leq\ell,j\leq n}]\leavevmode\nobreak\ ,$
where $f$ is the function introduced in Equation 4.1. This the determinant of the Gram matrix of $\underline{\lambda}$ with respect to the Hermitian pairing $f$.
@@
$f^{g}(\underline{\lambda})=g!\cdot f(\lambda_{1})\wedge\ldots\wedge f(\lambda_{g})\leavevmode\nobreak\ .$
@@
    $f(\lambda)\wedge f(\lambda)=0\leavevmode\nobreak\ ,$
@@
      $f(\lambda_{1},\lambda_{2})\wedge f(\lambda_{1},\lambda_{2})=0\leavevmode\nobreak\ ;$
@@
      $f(\lambda_{1},\lambda_{2})\wedge f(\lambda_{2},\lambda_{1})=-f(\lambda_{1})\wedge f(\lambda_{2})\leavevmode\nobreak\ ;$
@@
      $f(\lambda_{1})\wedge f(\lambda_{1},\lambda_{2})=0\leavevmode\nobreak\ .$
@@
    $f(u\lambda_{1}+v\lambda_{2})\wedge f(s\lambda_{1}+t\lambda_{2})=|\det(A)|^{2}f(\lambda_{1})\wedge f(\lambda_{2})\leavevmode\nobreak\ .$
@@
By 4.7, we can write $f(\lambda)=\frac{i}{\sqrt{d_{k}}}u_{\lambda}^{*}(dz)\wedge u_{\lambda}^{*}(d\overline{z})$, then:
@@
$=0\leavevmode\nobreak\ .$
This proves $(1)$. For $(2)-(a)$, we have:
@@
$=-f(\lambda_{1})\wedge f(\lambda_{2})\leavevmode\nobreak\ .$
We similarly prove $(2)-(c)$. For $(3)$, we can write by sesquilinearity of $f$:
@@
$=$ $|ut-vs|^{2}f(\lambda_{1})\wedge f(\lambda_{2})\leavevmode\nobreak\ .$
@@
Let $r\geq 1$ and let $\lambda_{1},\ldots,\lambda_{r}$ be vectors in $M_{\mathbb{R}}$. Then the following identity holds:
$\prod_{\ell=1}^{r}f(\lambda_{\ell},\lambda_{\ell+1})=(-1)^{r-1}f(\lambda_{1})\wedge\ldots\wedge f(\lambda_{k})$
where $\lambda_{r+1}=\lambda_{1}$.
@@
$\prod_{\ell=1}^{r+1}f(\lambda_{\ell},\lambda_{\ell+1})=\left(\frac{i}{\sqrt{d_{k}}}\right)^{r+1}\prod_{\ell=1}^{r+1}u_{\lambda_{\ell}}^{*}(dz)\wedge u_{\lambda_{\ell+1}}^{*}(d\overline{z})\\ =\left(\frac{i}{\sqrt{d_{k}}}\right)^{r+1}\left[\prod_{\ell=1}^{r-1}u_{\lambda_{\ell}}^{*}(dz)\wedge u_{\lambda_{\ell+1}}^{*}(d\overline{z})\right]\wedge u_{\lambda_{r}}^{*}(dz)\wedge u_{\lambda_{r+1}}^{*}(d\overline{z})\wedge u_{\lambda_{r+1}}^{*}(dz)\wedge u_{\lambda_{1}}^{*}(d\overline{z})\\ =-\left(\frac{i}{\sqrt{d_{k}}}\right)^{r+1}\left[\prod_{\ell=1}^{r-1}u_{\lambda_{\ell}}^{*}(dz)\wedge u_{\lambda_{\ell+1}}^{*}(d\overline{z})\right]\wedge u_{\lambda_{r}}^{*}(dz)\wedge u_{\lambda_{1}}^{*}(d\overline{z})\wedge u_{\lambda_{r+1}}^{*}(dz)\wedge u_{\lambda_{r+1}}^{*}(d\overline{z})\\ =(-1)^{r+1}f(\lambda_{1})\wedge\cdots\wedge f(\lambda_{r+1})\leavevmode\nobreak\ ,$
where we used the induction hypothesis in the last line. This concludes the proof.
We are now ready to prove Proposition 4.9.
Proof of Proposition 4.9.

### v2

@@ 4.1. Poincaré dual forms
For each $\lambda\in M^{\vee}$, let $u^{*}_{\lambda}(\omega_{E})$ denote the pullback of $\omega_{E}$ under the morphism in Equation˜3.1. Then $u^{*}_{\lambda}(\omega_{E})\in\mathcal{H}^{1,1}(E_{M},\mathbb{R})$, where $\mathcal{H}^{1,1}(E_{M},\mathbb{R})$ is the real vector space of harmonic $(1,1)$ forms on $E_{M}$ with real coefficients.
@@
$\lambda$ $\mapsto f(\lambda):=u^{*}_{\lambda}(\omega_{E})\penalty 10000\ .$
@@
$Z(\lambda_{1},\ldots,\lambda_{g})=Z(\lambda_{1})\cap\ldots\cap Z(\lambda_{g})\penalty 10000\ .$
As for $(3)$, it results from the equality $a]^{*}\omega_{E}=|a|^{2}\omega_{E}$ valid for $a\in\mathcal{O}_{k}$ and the composition:
$u_{a\lambda}:E_{M}\xrightarrow{h(\lambda,\,\cdot)}E\xrightarrow{[a]}E\penalty 10000\ .$
@@
$(x_{1},\ldots,x_{n})$ $\mapsto\sum_{\ell=1}^{n}h(\lambda,e_{\ell})\cdot x_{\ell}=\sum_{\ell=1}^{n}a_{\ell}\lambda_{\ell}\cdot x_{\ell}\$
@@
$Y_{\ell}]=a_{\ell}^{2}[E^{\ell-1}\times\{0\}\times E^{n-\ell+1}]\in\mathrm{H}^{2}(E_{M},\mathbb{Z})\penalty 10000\ .$
@@
$M^{\vee}=\oplus_{\ell=1}^{n}{\mathfrak{D}}^{-1}_{k}\frac{1}{a_{\ell}}e_{\ell}\subset M\otimes_{{\mathcal{O}}_{k}}k\penalty 10000\ .$
@@
$Z(\lambda)]=\sum_{\ell=1}^{n}|\lambda_{\ell}|^{2}Y_{\ell}-\sum_{1\leq\ell<j\leq n}\left[\Re(\lambda_{\ell}\overline{\lambda}_{j})Y_{\ell,j}^{+}+\frac{1}{\delta_{k}}\Im(\lambda_{\ell}\overline{\lambda}_{j})Y_{\ell j}^{-}\right]\in\mathrm{H}^{2}(E_{M},\mathbb{Q})\penalty 10000\ .$
@@
$u_{\lambda}^{*}(dz)=\sum_{\ell=1}^{n}a_{\ell}{\lambda}_{\ell}dz_{\ell}\quad\textrm{and}\quad u_{\lambda}^{*}(d\overline{z})=\sum_{\ell=1}^{n}a_{\ell}\overline{\lambda}_{\ell}d\overline{z}_{\ell}\penalty 10000\ .$
@@
$Y_{\ell,j}^{+}=\frac{ia_{\ell}a_{j}}{\sqrt{d_{k}}}(dz_{\ell}\wedge d\overline{z}_{j}+dz_{j}\wedge d\overline{z}_{\ell}),\quad\textrm{and}\quad Y_{\ell,j}^{-}=\delta_{k}\frac{ia_{\ell}a_{j}}{\sqrt{d_{k}}}(dz_{\ell}\wedge d\overline{z}_{j}-dz_{j}\wedge d\overline{z}_{\ell})\penalty 10000\ .$
@@
$Z(\lambda)]=\sum_{\ell=1}^{n}|\lambda_{\ell}|^{2}Y_{\ell}-\sum_{\ell<j}a_{\ell}a_{j}\Big(\Re(\lambda_{\ell}\overline{\lambda}_{j})Y_{\ell,j}^{+}+\Im(\lambda_{\ell}\overline{\lambda}_{j})Y^{-}_{\ell,j}\Big)\penalty 10000\ .$
@@
The class of the divisor $D_{{\mathfrak{J}}}$ satisfies:
$D_{{\mathfrak{J}}}=\sum_{\ell=1}^{n}\frac{1}{a_{\ell}}\cdot Y_{\ell}\in{H}^{2}(E_{M},\mathbb{Z})$
@@
The class of $D_{{\mathfrak{J}}}$ is determined by the Hermitian pairing $h$ on $\mathrm{H}_{1}(E_{M},\mathbb{C})=\mathrm{H}_{1}(E,\mathbb{C})\otimes M=T_{0}E\otimes M\oplus\overline{T_{0}E}\otimes M$. Notice that $D_{\mathfrak{J}}(\partial_{z}\otimes e_{\ell},\partial_{z}\otimes e_{j})=0$ and $D_{\mathfrak{J}}(\partial_{z}\otimes e_{\ell},\partial_{\overline{z}}\otimes e_{j})=\delta_{\ell j}\ \frac{i}{\sqrt{d_{k}}}\cdot a_{\ell}$. Therefore,
$D_{{\mathfrak{J}}}=\frac{i}{\sqrt{d_{k}}}\sum_{\ell=1}^{n}a_{\ell}\cdot dz_{\ell}\wedge d\overline{z}_{\ell}=\sum_{\ell=1}^{n}\frac{1}{a_{\ell}}\cdot Y_{\ell}\in{H}^{2}(E_{M},\mathbb{Z})\penalty 10000\ ,$
@@
We have the following relation between $D_{{\mathfrak{J}}}$ and the Chern class of the conormal bundle of ${\mathcal{B}}_{\mathfrak{J}}$:
$c_{1}({\mathcal{N}}_{{\mathcal{B}}_{{\mathfrak{J}}}}^{\vee})=\frac{d_{k}}{r_{\mathfrak{J}}}\cdot D_{{\mathfrak{J}}}\penalty 10000\ .$
@@
This follows from Theorem˜4.4 and Equation˜3.3. ∎
The following is a consequence of Proposition˜4.3, by unicity of harmonic lifts.
@@
$f(\lambda)=\sum_{\ell=1}^{n}|\lambda_{\ell}|^{2}f(e_{\ell})-\sum_{1\leq\ell<j\leq n}[\Re(\lambda_{\ell}\overline{\lambda}_{j})f_{\ell,j}^{+}+\frac{1}{\delta_{k}}\Im(\lambda_{\ell}\overline{\lambda}_{j})f_{\ell,j}^{-}]\in{\mathcal{H}}^{1,1}(E_{M},\mathbb{R})\penalty 10000\ .$
@@
$f(\lambda_{1},\lambda_{2})=\frac{i}{\sqrt{d_{k}}}u_{\lambda_{1}}^{*}(dz)\wedge u_{\lambda_{2}}^{*}(d\overline{z})\penalty 10000\ ,$
where $dz$ is the canonical harmonic $1$-form on $E=\mathbb{C}/{\mathcal{O}}_{k}$.
The next proposition shows that $f$ admits an extension to $M_{\mathbb{R}}$. We denote by $\Delta$ the $\mathrm{SU}(h)$-invariant Laplacian on $M_{\mathbb{R}}$.
@@
$f:M_{\mathbb{R}}\rightarrow\mathcal{H}^{1,1}(E_{M},\mathbb{C})\penalty 10000\ ,$
which is a Hermitian form and which satisfies $\Delta(f)=D_{\mathfrak{J}}$.
@@
It is clear from Theorem˜4.6 that $f$ admits an extension to $M_{\mathbb{R}}$. Moreover, the invariant Laplacian is given by:
$\Delta=\prescript{t}{}{\left(\frac{\partial}{\partial\lambda}\right)}\cdot H^{-1}\cdot\frac{\partial}{\partial\overline{\lambda}}=\sum_{\ell=1}^{n}\frac{1}{a_{\ell}}\frac{\partial^{2}}{\partial\lambda_{\ell}\partial\overline{\lambda}_{\ell}}\penalty 10000\ ,$
@@
$\Delta(f)(\lambda)=\sum_{\ell=1}^{n}\frac{1}{a_{\ell}}Y_{\ell},$
which proves the desired result by Theorem˜$4.$ ∎
@@
$\underline{\lambda}$ $\mapsto f(\underline{\lambda})=\det[\left(f(\lambda_{\ell},\lambda_{j})\right)_{1\leq\ell,j\leq n}]\penalty 10000\ ,$
where $f$ is the function introduced in Equation˜4.1. This the determinant of the Gram matrix of $\underline{\lambda}$ with respect to the Hermitian pairing $f$.
@@
$f^{g}(\underline{\lambda})=g!\cdot f(\lambda_{1})\wedge\ldots\wedge f(\lambda_{g})\penalty 10000\ .$
@@
    $f(\lambda)\wedge f(\lambda)=0\penalty 10000\ ,$
@@
      $f(\lambda_{1},\lambda_{2})\wedge f(\lambda_{1},\lambda_{2})=0\penalty 10000\ ;$
@@
      $f(\lambda_{1},\lambda_{2})\wedge f(\lambda_{2},\lambda_{1})=-f(\lambda_{1})\wedge f(\lambda_{2})\penalty 10000\ ;$
@@
      $f(\lambda_{1})\wedge f(\lambda_{1},\lambda_{2})=0\penalty 10000\ .$
@@
    $f(u\lambda_{1}+v\lambda_{2})\wedge f(s\lambda_{1}+t\lambda_{2})=|\det(A)|^{2}f(\lambda_{1})\wedge f(\lambda_{2})\penalty 10000\ .$
@@
By ˜4.7, we can write $f(\lambda)=\frac{i}{\sqrt{d_{k}}}u_{\lambda}^{*}(dz)\wedge u_{\lambda}^{*}(d\overline{z})$, then:
@@
$=0\penalty 10000\ .$
This proves $(1)$. For (2)(2)$(a)$, we have:
@@
$=-f(\lambda_{1})\wedge f(\lambda_{2})\penalty 10000\ .$
We similarly prove (2)(2)$(c)$. For $(3)$, we can write by sesquilinearity of $f$:
@@
$=$ $|ut-vs|^{2}f(\lambda_{1})\wedge f(\lambda_{2})\penalty 10000\ .$
@@
Let $r\geq 1$ and let $\lambda_{1},\ldots,\lambda_{r}$ be vectors in $M_{\mathbb{R}}$. Then the following identity holds $\prod_{\ell=1}^{r}f(\lambda_{\ell},\lambda_{\ell+1})=(-1)^{r-1}f(\lambda_{1})\wedge\ldots\wedge f(\lambda_{r})$, where $\lambda_{r+1}=\lambda_{1}$.
@@
$\prod_{\ell=1}^{r+1}f(\lambda_{\ell},\lambda_{\ell+1})=\left(\frac{i}{\sqrt{d_{k}}}\right)^{r+1}\prod_{\ell=1}^{r+1}u_{\lambda_{\ell}}^{*}(dz)\wedge u_{\lambda_{\ell+1}}^{*}(d\overline{z})\\ =\left(\frac{i}{\sqrt{d_{k}}}\right)^{r+1}\left[\prod_{\ell=1}^{r-1}u_{\lambda_{\ell}}^{*}(dz)\wedge u_{\lambda_{\ell+1}}^{*}(d\overline{z})\right]\wedge u_{\lambda_{r}}^{*}(dz)\wedge u_{\lambda_{r+1}}^{*}(d\overline{z})\wedge u_{\lambda_{r+1}}^{*}(dz)\wedge u_{\lambda_{1}}^{*}(d\overline{z})\\ =-\left(\frac{i}{\sqrt{d_{k}}}\right)^{r+1}\left[\prod_{\ell=1}^{r-1}u_{\lambda_{\ell}}^{*}(dz)\wedge u_{\lambda_{\ell+1}}^{*}(d\overline{z})\right]\wedge u_{\lambda_{r}}^{*}(dz)\wedge u_{\lambda_{1}}^{*}(d\overline{z})\wedge u_{\lambda_{r+1}}^{*}(dz)\wedge u_{\lambda_{r+1}}^{*}(d\overline{z})\\ =(-1)^{r+1}f(\lambda_{1})\wedge\cdots\wedge f(\lambda_{r+1})\penalty 10000\ ,$
where we used the induction hypothesis in the last line. This concludes the proof. ∎
We are now ready to prove Proposition˜4.9.
Proof of Proposition˜4.9.